4g NaOH is dissolved in 100ml water. How many ml of 2.5 M H2SO4 will be required for its complete neutralization?

Grade:12

2 Answers

Swapnil Saxena
12 years ago

4g = 0.1 mols of NaOH

Molarity = (0.1mols)/(0.1L)=1M

Accord to the equation

N1*V1=N2*V2

M1*V1=M2*V2*2(becoz H2SO4 is a diprotic acid)

1M*0.1=2.5*2*V2

=0.1/5 = 0.02L

Shyama P
12 years ago

20ml